OpenStudy (anonymous):

Please Help:( Solve each system by graphing. y=1/2x-2 y=7/4x=3 THANK YOU

OpenStudy (mysesshou):

\[y=(1/2)x-2\] \[y=(7/4)x-3\] assumed First one: Y intercept : (0,-2) X intercept: 0=1/2x-2 1/2x = 2 x=4 (4,0) Second one: Y intercept: (0,-3) X intercept: 0 = (7/4)x-3 7/4x=3 7x=12 x=12/7 ~1.7 (1.7,0)

OpenStudy (mysesshou):

 here's what I get. :)